Part – B:                                                                                                                              Marks [2]

Two points (p, q) on a linear supply function are ($2.55, 98,000) and ($6, 195,500).

  1. Determine the supply function q = f(p).
  2. What price would result in supplier offering 145,000 units for sale?
  3. Interpret the slope of the function.
  4. Determine the p intercept.
  5. Sketch f(p)

Part – C:                                                                                                                              Marks [2]

The supply and demand functions for a product are; qs=6p2-600 and qd=5p2-20p+900 . Determine the market equilibrium price and quantity. Also draw the supply-demand equilibrium graph.

Part – D:                                                                                                                              Marks [1]

A piece of machinery is purchased for $500,000. Accountants have decided to use a straight-line depreciation method with machine being fully depreciated after 12 years. Letting V equal the book value of the machine;

  1. Assuming that there is no salvage value, determine the function V=f (t).
  2. If the machine can be resold after 12 years for $60,000. Determine the function V=f (t).
Answer #1

Ans: As per Chegg answering guidelines, I can answer only 4 sub parts of the question.

(a) Determing supply function q = f(p).

  • q ≈ 28270p +25935
  • $4.21 will give a supply of 145,000
  • slope is the supply increase per dollar of price
  • p-intercept is about -$0.92

Step-by-step explanation:

The two-point form of the equation of a line is useful for this.

y = (y2 -y1)/(x2 -x1)(x -x1) +y1

q = (195500 -98000)/(6.00 -2.55)(p -2.55) +98000

q = 97500/3.45(p -2.55) +98000

q ≈ 28269.9p +25934.8

(b) Price that would result in supplier offering 145,000 units for sale

For a supply of 145000, the price would need to be ...

145000 = 28269.9p +25934.8

119065 = 28270p

119065/28270 = p = 4.21

A price of $4.21 will result in a supply of $145,000.

(c) Slope of the function

The slope is the additional amount supplied for each dollar increase in price.

(d) Determining the p-intercept

The price value that results in a supply of zero is ...

0 = 28270p +25935

-25935 = 28270p

-25935/28270 = p ≈ -0.92

The p-intercept is about -$0.92.